# Cislunar Constellation Design for Space Situational Awareness with Time-Expanded Facility Location Problem

### v1

@@ 7.2 LPOs used in TE-$p$-MP
Tables 10(d)11(d), and 12(d) provide a summary of LPOs used by solutions to the TE-$p$-MP by Gurobi with a time limit of 3600 seconds and the LM with a time limit of 1000 seconds.
Table 10: LPOs used in TE-$p$-MP instances with SOI demand from Gurobi and Lagrangean method solutions
@@
DPO (4:1) $\times$ 2
DPO (4:1) $\times$ 2
@@
L1 Lyapunov (1:1) $\times$ 2
Butterfly-S (1:1) $\times$ 1
Butterfly-N (1:1) $\times$ 1
@@
L1 Lyapunov (1:1) $\times$ 2
DRO (9:4) $\times$ 1
Butterfly-S (9:4) $\times$ 1
@@
DPO (4:1) $\times$ 3
DPO (4:1) $\times$ 3
@@
L1 Lyapunov (1:1) $\times$ 2
DPO (1:1) $\times$ 1
Butterfly-S (1:1) $\times$ 3
@@
L1 Lyapunov (1:1) $\times$ 1
DRO (3:2) $\times$ 2
Butterfly-N (9:4) $\times$ 2
Butterfly-S (9:4) $\times$ 1
@@
DPO (4:1) $\times$ 4
DPO (4:1) $\times$ 4
@@
DRO (5:2) $\times$ 1
L1 Lyapunov (1:1) $\times$ 1
DPO (1:1) $\times$ 1
L1 Lyapunov (2:1) $\times$ 1
Butterfly-S (1:1) $\times$ 2
Butterfly-N (1:1) $\times$ 2
@@
DRO (3:2) $\times$ 4
Butterfly-N (9:4) $\times$ 2
DRO (9:4) $\times$ 1
Butterfly-S (9:4) $\times$ 1
@@
DPO (3:1) $\times$ 2
Butterfly-N (2:1) $\times$ 2
Butterfly-S (2:1) $\times$ 1
DPO (4:1) $\times$ 4
DPO (3:1) $\times$ 1
@@
DRO (5:2) $\times$ 1
L1 Lyapunov (1:1) $\times$ 1
DPO (1:1) $\times$ 1
Butterfly-S (3:2) $\times$ 1
L1 Lyapunov (2:1) $\times$ 1
L1 Lyapunov (1:1) $\times$ 1
Butterfly-S (1:1) $\times$ 1
Butterfly-N (1:1) $\times$ 3
@@
DRO (3:2) $\times$ 5
L2 Halo (N) (9:2) $\times$ 1
Butterfly-N (9:4) $\times$ 1
L2 Halo (S) (9:2) $\times$ 2
Butterfly-S (9:4) $\times$ 1
Table 11: LPOs used in TE-$p$-MP instances with Cone of Shame demand from Gurobi and Lagrangean method solutions
@@
L2 Halo (N) (9:2) $\times$ 1
L2 Halo (S) (9:2) $\times$ 1
Butterfly-S (1:1) $\times$ 1
Butterfly-N (1:1) $\times$ 1
@@
L1 Lyapunov (1:1) $\times$ 2
L1 Lyapunov (1:1) $\times$ 2
@@
DPO (1:1) $\times$ 1
L1 Lyapunov (2:1) $\times$ 1
L1 Lyapunov (1:1) $\times$ 2
@@
L2 Halo (N) (9:2) $\times$ 1
L2 Halo (S) (9:2) $\times$ 1
Butterfly-N (1:1) $\times$ 1
Butterfly-N (1:1) $\times$ 3
@@
L1 Lyapunov (1:1) $\times$ 3
L1 Lyapunov (1:1) $\times$ 1
L2 Lyapunov (1:1) $\times$ 2
@@
DRO (5:2) $\times$ 1
DPO (1:1) $\times$ 1
L1 Lyapunov (2:1) $\times$ 1
L1 Lyapunov (1:1) $\times$ 3
@@
Butterfly-S (1:1) $\times$ 2
Butterfly-N (1:1) $\times$ 2
Butterfly-S (1:1) $\times$ 2
Butterfly-N (1:1) $\times$ 2
@@
L1 Lyapunov (1:1) $\times$ 3
L2 Lyapunov (1:1) $\times$ 1
L1 Lyapunov (1:1) $\times$ 1
L2 Lyapunov (1:1) $\times$ 3
@@
DRO (5:2) $\times$ 1
L1 Lyapunov (1:1) $\times$ 1
DPO (1:1) $\times$ 1
L1 Lyapunov (2:1) $\times$ 1
L1 Lyapunov (1:1) $\times$ 4
@@
L2 Halo (N) (9:2) $\times$ 1
Butterfly-S (1:1) $\times$ 1
Butterfly-S (3:2) $\times$ 1
L2 Halo (S) (9:2) $\times$ 1
Butterfly-N (1:1) $\times$ 1
Butterfly-S (1:1) $\times$ 4
Butterfly-N (1:1) $\times$ 1
@@
L1 Lyapunov (1:1) $\times$ 2
Butterfly-S (1:1) $\times$ 1
L2 Lyapunov (1:1) $\times$ 2
L2 Lyapunov (1:1) $\times$ 5
@@
DRO (5:2) $\times$ 1
L1 Lyapunov (1:1) $\times$ 1
DPO (1:1) $\times$ 1
Butterfly-S (3:2) $\times$ 1
L1 Lyapunov (2:1) $\times$ 1
L1 Lyapunov (1:1) $\times$ 5
Table 12: LPOs used in TE-$p$-MP instances with LET demand from Gurobi and Lagrangean method solutions
(a) $p=2$
[…]
15
L2 Halo (N) (2:1) $\times$ 1
L2 Halo (S) (2:1) $\times$ 1
L2 Halo (S) (2:1) $\times$ 2
18
L1 Lyapunov (1:1) $\times$ 2
L2 Halo (N) (3:1) $\times$ 2
20
L1 Lyapunov (1:1) $\times$ 2
L2 Halo (N) (3:1) $\times$ 2
(b) $p=3$
[…]
15
L2 Halo (N) (2:1) $\times$ 1
L2 Halo (S) (2:1) $\times$ 2
L2 Halo (S) (2:1) $\times$ 3
18
DRO (3:2) $\times$ 1
L2 Lyapunov (1:1) $\times$ 2
L2 Halo (N) (3:1) $\times$ 3
20
DRO (5:2) $\times$ 1
DPO (1:1) $\times$ 1
L1 Lyapunov (2:1) $\times$ 1
L2 Halo (N) (3:1) $\times$ 3
(c) $p=4$
[…]
15
L2 Halo (N) (2:1) $\times$ 2
L2 Halo (S) (2:1) $\times$ 2
L2 Halo (S) (2:1) $\times$ 4
18
DRO (5:2) $\times$ 1
L1 Lyapunov (1:1) $\times$ 1
DPO (1:1) $\times$ 1
L1 Lyapunov (2:1) $\times$ 1
L2 Halo (N) (3:1) $\times$ 4
20
DRO (3:2) $\times$ 4
L2 Halo (N) (3:1) $\times$ 4
(d) $p=5$
[…]
15
L2 Halo (N) (2:1) $\times$ 2
L2 Halo (S) (2:1) $\times$ 3
L2 Halo (N) (2:1) $\times$ 1
L2 Halo (S) (2:1) $\times$ 4
18
DRO (2:1) $\times$ 1
DRO (3:2) $\times$ 4
L2 Halo (N) (3:1) $\times$ 5
20
DRO (3:2) $\times$ 5
L1 Lyapunov (1:1) $\times$ 1
L2 Halo (N) (3:1) $\times$ 4

### v2

@@ 7.2 LPOs used in TE-$p$-MP
Tables  8(d) and 9(d) provide a summary of LPOs used by solutions to the TE-$p$-MP by B&B with a time limit of 3600 seconds and the LM with a time limit of 500 seconds.
Table 8: LPOs used in TE-$p$-MP instances with Cone of Shame demand from Branch-and-Bound and Lagrangian method solutions
@@
halo L2N (9:2) $\times$ 1
halo L2S (9:2) $\times$ 1
butterflyS (1:1) $\times$ 1
butterflyN (1:1) $\times$ 1
@@
lyapunov L1 (1:1) $\times$ 2
lyapunov L1 (1:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
@@
lyapunov L2 (3:2) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
lyapunov L1 (1:1) $\times$ 1
halo L2N (3:1) $\times$ 1
@@
halo L2N (9:2) $\times$ 1
butterflyS (1:1) $\times$ 1
halo L2S (9:2) $\times$ 1
butterflyS (1:1) $\times$ 1
butterflyN (1:1) $\times$ 2
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
lyapunov L1 (1:1) $\times$ 3
@@
lyapunov L2 (1:1) $\times$ 3
lyapunov L1 (1:1) $\times$ 3
@@
halo L2N (9:2) $\times$ 1
butterflyS (1:1) $\times$ 1
halo L2S (9:2) $\times$ 1
butterflyN (1:1) $\times$ 1
butterflyS (1:1) $\times$ 2
butterflyN (1:1) $\times$ 2
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
dpo (9:4) $\times$ 1
lyapunov L1 (1:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 3
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
dpo (9:4) $\times$ 1
lyapunov L1 (1:1) $\times$ 4
@@
halo L2N (4:1) $\times$ 1
butterflyS (1:1) $\times$ 1
butterflyN (3:2) $\times$ 1
butterflyN (1:1) $\times$ 2
lyapunov L1 (9:4) $\times$ 1
butterflyN (1:1) $\times$ 4
@@
halo L2N (9:4) $\times$ 1
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
dpo (9:4) $\times$ 1
lyapunov L1 (1:1) $\times$ 3
lyapunov L2 (1:1) $\times$ 2
@@
lyapunov L1 (1:1) $\times$ 3
lyapunov L1 (3:2) $\times$ 1
dro (2:1) $\times$ 1
lyapunov L1 (1:1) $\times$ 5
Table 9: LPOs used in TE-$p$-MP instances with LET demand from Branch-and-Bound and Lagrangian method solutions
@@
halo L2N (2:1) $\times$ 1
halo L2S (2:1) $\times$ 1
halo L2S (2:1) $\times$ 2
@@
lyapunov L1 (1:1) $\times$ 2
halo L2N (3:1) $\times$ 2
@@
lyapunov L1 (1:1) $\times$ 2
halo L2N (3:1) $\times$ 2
@@
halo L2N (2:1) $\times$ 1
halo L2S (2:1) $\times$ 2
halo L2S (2:1) $\times$ 3
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
halo L2N (3:1) $\times$ 3
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
halo L2N (3:1) $\times$ 3
@@
halo L2N (2:1) $\times$ 2
halo L2S (2:1) $\times$ 2
halo L2S (2:1) $\times$ 4
@@
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
dpo (9:4) $\times$ 1
halo L2N (3:1) $\times$ 4
@@
dro (3:2) $\times$ 4
halo L2N (3:1) $\times$ 4
@@
halo L2N (2:1) $\times$ 2
halo L2S (2:1) $\times$ 3
halo L2N (2:1) $\times$ 1
halo L2S (2:1) $\times$ 4
@@
halo L2N (9:4) $\times$ 1
lyapunov L2 (3:2) $\times$ 1
butterflyN (2:1) $\times$ 1
lyapunov L2 (1:1) $\times$ 1
dpo (9:4) $\times$ 1
halo L2N (3:1) $\times$ 5
@@
halo L2N (3:1) $\times$ 1
dro (9:4) $\times$ 1
dro (3:2) $\times$ 3
lyapunov L1 (1:1) $\times$ 1
halo L2N (3:1) $\times$ 4